OpenStudy (anonymous):

Find the eighth term of the sequence given by the rule . 31 49 45 30

OpenStudy (anonymous):

Please help

OpenStudy (anonymous):

@StudyGurl14 @mathmath333

OpenStudy (studygurl14):

What's the rule?

OpenStudy (anonymous):

t^n=5n-9

OpenStudy (anonymous):

Big t little n sorry

OpenStudy (studygurl14):

\(\large t_n=5n-9\) Like this?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

@StudyGurl14 yes

OpenStudy (studygurl14):

"eighth term" Just plug in 8 for n and solve`

OpenStudy (studygurl14):

\(\large t_8=5(8)-9\)

OpenStudy (anonymous):

ah okay

OpenStudy (anonymous):

can u help me with 1 more ? @StudyGurl14 ?

OpenStudy (studygurl14):

sure. but i have to leave soon

OpenStudy (studygurl14):

don't post it here

OpenStudy (anonymous):

Find the third term of the sequence given by the rule . the rule is Tn=2(3)n-1 54 18 27 9

OpenStudy (studygurl14):

post it in a new question so other people can see it

OpenStudy (studygurl14):

\(\large T_n=2(3)^{n-1}\) Just plug in 3 for n and solve again

OpenStudy (studygurl14):

gtg bye

OpenStudy (anonymous):

okay